# Beyond the Kolmogorov Barrier: A Learnable Weighted Hybrid Autoencoder for Model Order Reduction

### v3

@@ 5 Conclusions
In this work, we present a novel deep autoencoder framework that demonstrates convergence properties akin to SVD. By incorporating a learnable weighted average between SVD and vanilla deep autoencoders (either feedforward or convolutional), our approach achieves SVD-like convergence as the rank increases. We validate the effectiveness of this framework on pure reconstruction tasks using two challenging chaotic PDE datasets: the 1D Kuramoto-Sivashinsky and the 3D homogeneous isotropic turbulence. The results show that our learnable weighted hybrid autoencoder consistently achieves the lowest testing error and exhibits superior robustness to noisy data compared to other methods such as POD, vanilla deep autoencoders, and simple hybrid autoencoders. Remarkably, we find that our proposed approach leads to a minimum with a sharpness that is a thousand times smaller than that of other deep autoencoder frameworks. In addition, we demonstrate that utilizing our proposed framework in tandem with time series prediction models, we can achieve superior performance for surrogate modeling of time-dependent PDE over other approaches. Our framework is also capable of capturing dynamics of the system with strong discontinuities without spurious oscillations in the solution. This highlights its potential for robust and generalizable representation learning in complex PDE systems.
@@
Acknowledgements
This work was supported by U.S. Department of Energy under Advancements in Artificial Intelligence for Science with award ID DE-SC0025425. The authors thank the Center for Computational Innovations (CCI) at Rensselaer Polytechnic Institute (RPI) for providing computational resources during the early stages of this research. Numerical experiments are performed using computational resources granted by NSF-ACCESS for the project PHY240112 and that of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using the NERSC award NERSC DDR-ERCAP0030714.
@@ Appendix A Detailed comparison of numerical results
We summarize the detailed results of all numerical experiments conducted in this work in tables 5 and 6. Upon publication, the code will be available at https://github.com/csml-rpi/deep-ae-with-svd-convergence.
Grid Rank Mean Train Error, $\mu$(±$\pm\sigma$) Mean Test Error, $\mu$(±$\pm\sigma$) Total Number of Model Parameters
@@
512 50 2.84e-01 2.54e-03 2.41e-04 1.62e-04 3.06e-01 2.73e-03 2.55e-04 1.92e-04 25,600 113,162 138,762 138,813
($\pm$ 1.92e-03) ($\pm$1.42e-03) ($\pm$6.07e-05) ($\pm$9.47e-06) ($\pm$4.80e-03) ($\pm$1.49e-03) ($\pm$5.96e-05) ($\pm$1.03e-05)
60 6.89e-02 1.68e-03 1.88e-04 1.59e-05 7.50e-02 1.76e-03 1.94e-04 1.96e-05 30,720 138,092 168,812 168,873
($\pm$ 4.92e-04) ($\pm$5.84e-04) ($\pm$7.35e-05) ($\pm$7.32e-07) ($\pm$1.47e-03) ($\pm$5.75e-04) ($\pm$7.38e-05) ($\pm$2.94e-06)
70 1.83e-02 1.75e-03 1.93e-04 2.73e-06 2.03e-02 1.79e-03 1.97e-04 3.55e-06 35,840 163,822 199,662 199,733
($\pm$ 1.37e-04) ($\pm$1.14e-03) ($\pm$1.26e-04) ($\pm$3.95e-07) ($\pm$3.67e-04) ($\pm$1.10e-03) ($\pm$1.28e-04) ($\pm$9.03e-07)
80 4.59e-03 2.07e-03 1.55e-04 1.02e-06 5.16e-03 2.22e-03 1.58e-04 1.39e-06 40,960 190,352 231,312 231,393
($\pm$ 2.95e-05) ($\pm$1.68e-03) ($\pm$4.62e-05) ($\pm$1.05e-07) ($\pm$8.78e-05) ($\pm$1.75e-03) ($\pm$4.72e-05) ($\pm$5.40e-07)
1024 50 3.18e-01 6.48e-03 3.01e-04 2.53e-04 3.41e-01 6.95e-03 3.20e-04 2.98e-04 51,200 216,074 267,274 267,325
($\pm$ 2.28e-03) ($\pm$2.53e-03) ($\pm$6.88e-05) ($\pm$2.29e-05) ($\pm$5.12e-03) ($\pm$2.49e-03) ($\pm$6.89e-05) ($\pm$2.29e-05)
60 8.75e-02 6.10e-03 2.14e-04 2.89e-05 9.53e-02 6.19e-03 2.22e-04 3.61e-05 61,440 261,484 322,924 322,985
($\pm$ 4.46e-04) ($\pm$7.01e-03) ($\pm$6.56e-05) ($\pm$3.29e-06) ($\pm$1.11e-03) ($\pm$6.43e-03) ($\pm$6.54e-05) ($\pm$5.84e-06)
70 2.21e-02 3.76e-03 2.25e-04 5.33e-06 2.46e-02 4.08e-03 2.30e-04 7.10e-06 71,680 307,694 379,374 379,445
($\pm$ 9.31e-05) ($\pm$2.64e-03) ($\pm$1.31e-04) ($\pm$1.31e-06) ($\pm$3.32e-04) ($\pm$2.72e-03) ($\pm$1.30e-04) ($\pm$2.38e-06)
80 5.86e-03 3.69e-03 2.23e-04 1.94e-06 6.58e-03 4.00e-03 2.26e-04 2.68e-06 81,920 354,704 436,624 436,705
($\pm$ 3.16e-05) ($\pm$2.34e-03) ($\pm$1.07e-04) ($\pm$1.54e-07) ($\pm$9.69e-05) ($\pm$2.34e-03) ($\pm$1.06e-04) ($\pm$8.49e-07)
2048 50 2.92e-01 1.62e-02 2.74e-04 2.36e-04 3.15e-01 1.82e-02 2.98e-04 2.80e-04 102,400 421,898 524,298 524,349
($\pm$ 2.34e-03) ($\pm$1.81e-02) ($\pm$7.71e-05) ($\pm$2.03e-05) ($\pm$5.32e-03) ($\pm$1.88e-02) ($\pm$8.97e-05) ($\pm$3.50e-05)
60 7.20e-02 9.93e-03 1.65e-04 3.39e-05 7.84e-02 1.14e-02 1.76e-04 4.35e-05 122,880 508,268 631,148 631,209
($\pm$ 4.95e-04) ($\pm$4.29e-03) ($\pm$2.48e-05) ($\pm$2.63e-06) ($\pm$1.31e-03) ($\pm$4.79e-03) ($\pm$2.31e-05) ($\pm$1.05e-05)
70 1.80e-02 7.65e-03 2.52e-04 1.21e-05 1.99e-02 9.18e-03 2.59e-04 1.61e-05 143,360 595,438 738,798 738,869
($\pm$ 9.85e-05) ($\pm$3.15e-03) ($\pm$7.35e-05) ($\pm$6.28e-07) ($\pm$3.20e-04) ($\pm$3.47e-03) ($\pm$7.46e-05) ($\pm$4.49e-06)
80 4.72e-03 9.62e-03 3.07e-04 9.23e-06 5.34e-03 1.12e-02 3.13e-04 1.23e-05 163,840 683,408 847,248 847,329
($\pm$ 3.99e-05) ($\pm$6.61e-03) ($\pm$2.34e-04) ($\pm$3.11e-07) ($\pm$1.54e-04) ($\pm$6.77e-03) ($\pm$2.30e-04) ($\pm$2.95e-06)
@@
Table 5: Summary of training $L^{2}$ error, testing $L^{2}$ error, and the number of parameters for the four models trained on data with varying grid resolutions and ranks $r$ pertaining to 1D K-S case. The standard deviation of the error is indicated in parenthesis. The number of parameters for the hybrid approaches includes non-trainable POD parameters, which remain fixed throughout the optimization process and are not trainable.
@@
Grid Rank Mean Train Error, $\mu$(±$\pm\sigma$) Mean Test Error, $\mu$(±$\pm\sigma$) Total Number of Model Parameters
@@
$16^{3}$ 5 2.01e+01 1.05e+00 1.35e+00 1.76e+00 2.04e+01 5.74e+00 5.98e+00 5.48e+00 61,440 688,348,168 688,409,608 688,409,616
($\pm$ 0.00) ($\pm$1.37e-01) ($\pm$1.13e-01) ($\pm$9.31e-02) ($\pm$0.00) ($\pm$3.13e-01) ($\pm$1.09e-01) ($\pm$6.44e-02)
10 1.41e+01 9.71e-01 6.85e-01 9.02e-01 1.47e+01 5.53e+00 5.17e+00 5.22e+00 122,880 688,368,653 688,491,533 688,491,546
($\pm$ 0.0) ($\pm$2.75e-01) ($\pm$6.05e-02) ($\pm$4.04e-02) ($\pm$0.0) ($\pm$1.11e-01) ($\pm$3.20e-02) ($\pm$3.49e-02)
15 1.14e+01 8.76e-01 4.65e-01 6.35e-01 1.23e+01 5.47e+00 5.08e+00 5.20e+00 184,320 688,389,138 688,573,458 688,573,476
($\pm$ 0.0) ($\pm$4.36e-02) ($\pm$2.43e-02) ($\pm$3.55e-02) ($\pm$0.00) ($\pm$1.02e-01) ($\pm$3.22e-02) ($\pm$3.59e-02)
20 9.71e+00 8.96e-01 3.07e-01 5.06e-01 1.09e+01 5.49e+00 4.85e+00 5.21e+00 245,760 688,409,623 688,655,383 688,655,406
($\pm$ 0.00) ($\pm$5.07e-02) ($\pm$1.90e-02) ($\pm$2.83e-02) ($\pm$0.00) ($\pm$6.92e-02) ($\pm$3.01e-02) ($\pm$3.14e-02)
$32^{3}$ 5 2.02e+01 9.15e+00 1.84e+01 5.56e+00 2.04e+01 1.05e+01 1.87e+01 7.16e+00 491,520 688,505,864 688,997,384 688,997,392
($\pm$ 0.00) ($\pm$1.10e+01) ($\pm$2.61e-01) ($\pm$1.03e-01) ($\pm$0.00) ($\pm$1.05e+01) ($\pm$2.34e-01) ($\pm$7.28e-02)
10 1.42e+01 1.52e+01 1.23e+01 3.98e+00 1.47e+01 1.63e+01 1.32e+01 6.50e+00 983,040 688,669,709 689,652,749 689,652,762
($\pm$ 0.00) ($\pm$1.74e+01) ($\pm$1.45e-01) ($\pm$6.67e-02) ($\pm$0.00) ($\pm$1.66e+01) ($\pm$1.27e-01) ($\pm$5.38e-02)
15 1.15e+01 1.46e+01 9.18e+00 2.90e+00 1.22e+01 1.59e+01 1.06e+01 6.18e+00 1,474,560 688,833,554 690,308,114 690,308,132
($\pm$ 0.00) ($\pm$1.76e+01) ($\pm$1.81e-01) ($\pm$5.54e-02) ($\pm$0.00) ($\pm$1.68e+01) ($\pm$1.55e-01) ($\pm$4.65e-02)
20 9.78e+00 1.10e+01 6.91e+00 2.21e+00 1.08e+01 1.26e+01 8.86e+00 5.97e+00 1,966,080 688,997,399 690,963,479 690,963,502
($\pm$ 0.00) ($\pm$1.52e+01) ($\pm$2.84e-01) ($\pm$4.70e-02) ($\pm$0.00) ($\pm$1.44e+01) ($\pm$1.55e-01) ($\pm$3.96e-02)
$64^{3}$ 5 2.02e+01 2.89e+01 1.95e+01 9.00e+00 2.04e+01 2.89e+01 1.96e+01 9.53e+00 3,932,160 689,767,432 693,699,592 693,699,600
($\pm$ 0.00) ($\pm$1.94e+01) ($\pm$1.53e-01) ($\pm$9.62e-02) ($\pm$0.00) ($\pm$1.91e+01) ($\pm$1.48e-01) ($\pm$9.27e-02)
10 1.42e+01 2.50e+01 1.35e+01 7.01e+00 1.46e+01 2.52e+01 1.40e+01 7.95e+00 7,864,320 691,078,157 698,942,477 698,942,490
($\pm$ 0.00) ($\pm$2.00e+01) ($\pm$1.22e-01) ($\pm$7.07e-02) ($\pm$0.00) ($\pm$1.96e+01) ($\pm$1.18e-01) ($\pm$6.18e-02)
15 1.15e+01 1.65e+01 1.08e+01 5.59e+00 1.22e+01 1.69e+01 1.16e+01 7.01e+00 11,796,480 692,388,882 704,185,362 704,185,380
($\pm$ 0.00) ($\pm$1.65e+01) ($\pm$1.79e-01) ($\pm$3.82e-02) ($\pm$0.00) ($\pm$1.62e+01) ($\pm$1.69e-01) ($\pm$3.63e-02)
20 9.78e+00 1.94e+01 9.13e+00 4.54e+00 1.08e+01 1.97e+01 1.02e+01 6.45e+00 15,728,640 693,699,607 709,428,247 709,428,270
($\pm$ 0.00) ($\pm$1.82e+01) ($\pm$9.70e-02) ($\pm$3.73e-02) ($\pm$0.00) ($\pm$1.78e+01) ($\pm$9.29e-02) ($\pm$3.39e-02)

### v4

@@ 5 Conclusions
In this work, we present a novel deep autoencoder framework that demonstrates convergence properties akin to SVD. By incorporating a learnable weighted average between SVD and vanilla deep autoencoders (either feedforward or convolutional), our approach achieves SVD-like convergence as the rank increases. We validate the effectiveness of this framework on pure reconstruction tasks using two challenging chaotic PDE datasets: the 1D Kuramoto-Sivashinsky and the 3D homogeneous isotropic turbulence. The results show that our learnable weighted hybrid autoencoder consistently achieves the lowest testing error and exhibits superior robustness to noisy data compared to other methods such as POD, vanilla deep autoencoders, and simple hybrid autoencoders. Remarkably, we find that our proposed approach leads to a minimum with a sharpness that is a thousand times smaller than that of other deep autoencoder frameworks. In addition, we demonstrate that utilizing our proposed framework in tandem with time series prediction models, we can achieve superior performance for surrogate modeling of time-dependent PDEs over other approaches. Our framework is also capable of capturing dynamics of the system with strong discontinuities without spurious oscillations in the solution. This highlights its potential for robust and generalizable representation learning in complex PDE systems.
@@
Acknowledgement
This work was supported by U.S. Department of Energy under Advancements in Artificial Intelligence for Science with award number DE-SC0025425. The authors thank the Center for Computational Innovations (CCI) at Rensselaer Polytechnic Institute (RPI) for providing computational resources during the early stages of this research. Numerical experiments are performed using computational resources granted by NSF-ACCESS for the project PHY240112 and that of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using the NERSC award NERSC DDR-ERCAP0030714.
@@ Appendix A Detailed comparison of numerical results
We summarize the detailed results of all numerical experiments conducted in this work in Tables 6 and 7. Upon publication, the code and data will be available at https://github.com/csml-rpi/deep-ae-with-svd-convergence.
Grid Rank Mean Train Error, $\mu$($\pm\sigma$) Mean Test Error, $\mu$($\pm\sigma$) Total Number of Model Parameters
@@
512 50 $2.84\times 10^{-1}$ $2.54\times 10^{-3}$ $2.41\times 10^{-4}$ $1.62\times 10^{-4}$ $3.06\times 10^{-1}$ $2.73\times 10^{-3}$ $2.55\times 10^{-4}$ $1.92\times 10^{-4}$ 25,600 113,162 138,762 138,813
($\pm 1.92\times 10^{-3}$) ($\pm 1.42\times 10^{-3}$) ($\pm 6.07\times 10^{-5}$) ($\pm 9.47\times 10^{-6}$) ($\pm 4.80\times 10^{-3}$) ($\pm 1.49\times 10^{-3}$) ($\pm 5.96\times 10^{-5}$) ($\pm 1.03\times 10^{-5}$)
60 $6.89\times 10^{-2}$ $1.68\times 10^{-3}$ $1.88\times 10^{-4}$ $1.59\times 10^{-5}$ $7.50\times 10^{-2}$ $1.76\times 10^{-3}$ $1.94\times 10^{-4}$ $1.96\times 10^{-5}$ 30,720 138,092 168,812 168,873
($\pm 4.92\times 10^{-4}$) ($\pm 5.84\times 10^{-4}$) ($\pm 7.35\times 10^{-5}$) ($\pm 7.32\times 10^{-7}$) ($\pm 1.47\times 10^{-3}$) ($\pm 5.75\times 10^{-4}$) ($\pm 7.38\times 10^{-5}$) ($\pm 2.94\times 10^{-6}$)
70 $1.83\times 10^{-2}$ $1.75\times 10^{-3}$ $1.93\times 10^{-4}$ $2.73\times 10^{-6}$ $2.03\times 10^{-2}$ $1.79\times 10^{-3}$ $1.97\times 10^{-4}$ $3.55\times 10^{-6}$ 35,840 163,822 199,662 199,733
($\pm 1.37\times 10^{-4}$) ($\pm 1.14\times 10^{-3}$) ($\pm 1.26\times 10^{-4}$) ($\pm 3.95\times 10^{-7}$) ($\pm 3.67\times 10^{-4}$) ($\pm 1.10\times 10^{-3}$) ($\pm 1.28\times 10^{-4}$) ($\pm 9.03\times 10^{-7}$)
80 $4.59\times 10^{-3}$ $2.07\times 10^{-3}$ $1.55\times 10^{-4}$ $1.02\times 10^{-6}$ $5.16\times 10^{-3}$ $2.22\times 10^{-3}$ $1.58\times 10^{-4}$ $1.39\times 10^{-6}$ 40,960 190,352 231,312 231,393
($\pm 2.95\times 10^{-5}$) ($\pm 1.68\times 10^{-3}$) ($\pm 4.62\times 10^{-5}$) ($\pm 1.05\times 10^{-7}$) ($\pm 8.78\times 10^{-5}$) ($\pm 1.75\times 10^{-3}$) ($\pm 4.72\times 10^{-5}$) ($\pm 5.40\times 10^{-7}$)
1024 50 $3.18\times 10^{-1}$ $6.48\times 10^{-3}$ $3.01\times 10^{-4}$ $2.53\times 10^{-4}$ $3.41\times 10^{-1}$ $6.95\times 10^{-3}$ $3.20\times 10^{-4}$ $2.98\times 10^{-4}$ 51,200 216,074 267,274 267,325
($\pm 2.28\times 10^{-3}$) ($\pm 2.53\times 10^{-3}$) ($\pm 6.88\times 10^{-5}$) ($\pm 2.29\times 10^{-5}$) ($\pm 5.12\times 10^{-3}$) ($\pm 2.49\times 10^{-3}$) ($\pm 6.89\times 10^{-5}$) ($\pm 2.29\times 10^{-5}$)
60 $8.75\times 10^{-2}$ $6.10\times 10^{-3}$ $2.14\times 10^{-4}$ $2.89\times 10^{-5}$ $9.53\times 10^{-2}$ $6.19\times 10^{-3}$ $2.22\times 10^{-4}$ $3.61\times 10^{-5}$ 61,440 261,484 322,924 322,985
($\pm 4.46\times 10^{-4}$) ($\pm 7.01\times 10^{-3}$) ($\pm 6.56\times 10^{-5}$) ($\pm 3.29\times 10^{-6}$) ($\pm 1.11\times 10^{-3}$) ($\pm 6.43\times 10^{-3}$) ($\pm 6.54\times 10^{-5}$) ($\pm 5.84\times 10^{-6}$)
70 $2.21\times 10^{-2}$ $3.76\times 10^{-3}$ $2.25\times 10^{-4}$ $5.33\times 10^{-6}$ $2.46\times 10^{-2}$ $4.08\times 10^{-3}$ $2.30\times 10^{-4}$ $7.10\times 10^{-6}$ 71,680 307,694 379,374 379,445
($\pm 9.31\times 10^{-5}$) ($\pm 2.64\times 10^{-3}$) ($\pm 1.31\times 10^{-4}$) ($\pm 1.31\times 10^{-6}$) ($\pm 3.32\times 10^{-4}$) ($\pm 2.72\times 10^{-3}$) ($\pm 1.30\times 10^{-4}$) ($\pm 2.38\times 10^{-6}$)
80 $5.86\times 10^{-3}$ $3.69\times 10^{-3}$ $2.23\times 10^{-4}$ $1.94\times 10^{-6}$ $6.58\times 10^{-3}$ $4.00\times 10^{-3}$ $2.26\times 10^{-4}$ $2.68\times 10^{-6}$ 81,920 354,704 436,624 436,705
($\pm 3.16\times 10^{-5}$) ($\pm 2.34\times 10^{-3}$) ($\pm 1.07\times 10^{-4}$) ($\pm 1.54\times 10^{-7}$) ($\pm 9.69\times 10^{-5}$) ($\pm 2.34\times 10^{-3}$) ($\pm 1.06\times 10^{-4}$) ($\pm 8.49\times 10^{-7}$)
2048 50 $2.92\times 10^{-1}$ $1.62\times 10^{-2}$ $2.74\times 10^{-4}$ $2.36\times 10^{-4}$ $3.15\times 10^{-1}$ $1.82\times 10^{-2}$ $2.98\times 10^{-4}$ $2.80\times 10^{-4}$ 102,400 421,898 524,298 524,349
($\pm 2.34\times 10^{-3}$) ($\pm 1.81\times 10^{-2}$) ($\pm 7.71\times 10^{-5}$) ($\pm 2.03\times 10^{-5}$) ($\pm 5.32\times 10^{-3}$) ($\pm 1.88\times 10^{-2}$) ($\pm 8.97\times 10^{-5}$) ($\pm 3.50\times 10^{-5}$)
60 $7.20\times 10^{-2}$ $9.93\times 10^{-3}$ $1.65\times 10^{-4}$ $3.39\times 10^{-5}$ $7.84\times 10^{-2}$ $1.14\times 10^{-2}$ $1.76\times 10^{-4}$ $4.35\times 10^{-5}$ 122,880 508,268 631,148 631,209
($\pm 4.95\times 10^{-4}$) ($\pm 4.29\times 10^{-3}$) ($\pm 2.48\times 10^{-5}$) ($\pm 2.63\times 10^{-6}$) ($\pm 1.31\times 10^{-3}$) ($\pm 4.79\times 10^{-3}$) ($\pm 2.31\times 10^{-5}$) ($\pm 1.05\times 10^{-5}$)
70 $1.80\times 10^{-2}$ $7.65\times 10^{-3}$ $2.52\times 10^{-4}$ $1.21\times 10^{-5}$ $1.99\times 10^{-2}$ $9.18\times 10^{-3}$ $2.59\times 10^{-4}$ $1.61\times 10^{-5}$ 143,360 595,438 738,798 738,869
($\pm 9.85\times 10^{-5}$) ($\pm 3.15\times 10^{-3}$) ($\pm 7.35\times 10^{-5}$) ($\pm 6.28\times 10^{-7}$) ($\pm 3.20\times 10^{-4}$) ($\pm 3.47\times 10^{-3}$) ($\pm 7.46\times 10^{-5}$) ($\pm 4.49\times 10^{-6}$)
80 $4.72\times 10^{-3}$ $9.62\times 10^{-3}$ $3.07\times 10^{-4}$ $9.23\times 10^{-6}$ $5.34\times 10^{-3}$ $1.12\times 10^{-2}$ $3.13\times 10^{-4}$ $1.23\times 10^{-5}$ 163,840 683,408 847,248 847,329
($\pm 3.99\times 10^{-5}$) ($\pm 6.61\times 10^{-3}$) ($\pm 2.34\times 10^{-4}$) ($\pm 3.11\times 10^{-7}$) ($\pm 1.54\times 10^{-4}$) ($\pm 6.77\times 10^{-3}$) ($\pm 2.30\times 10^{-4}$) ($\pm 2.95\times 10^{-6}$)
@@
Table 6: Summary of training $L^{2}$ error, testing $L^{2}$ error, and the number of parameters for the four models trained on data with varying grid resolutions and ranks $r$ pertaining to 1D K-S case. The standard deviation of the error is indicated in parentheses. The number of parameters for the hybrid approaches includes non-trainable POD parameters, which remain fixed throughout the optimization process and are not trainable.
@@
Grid Rank Mean Train Error, $\mu$($\pm\sigma$) Mean Test Error, $\mu$($\pm\sigma$) Total Number of Model Parameters
@@
$16^{3}$ 5 $2.01\times 10^{1}$ $1.05\times 10^{0}$ $1.35\times 10^{0}$ $1.76\times 10^{0}$ $2.04\times 10^{1}$ $5.74\times 10^{0}$ $5.98\times 10^{0}$ $5.48\times 10^{0}$ 61,440 688,348,168 688,409,608 688,409,616
($\pm 0.00$) ($\pm 1.37\times 10^{-1}$) ($\pm 1.13\times 10^{-1}$) ($\pm 9.31\times 10^{-2}$) ($\pm 0.00$) ($\pm 3.13\times 10^{-1}$) ($\pm 1.09\times 10^{-1}$) ($\pm 6.44\times 10^{-2}$)
10 $1.41\times 10^{1}$ $9.71\times 10^{-1}$ $6.85\times 10^{-1}$ $9.02\times 10^{-1}$ $1.47\times 10^{1}$ $5.53\times 10^{0}$ $5.17\times 10^{0}$ $5.22\times 10^{0}$ 122,880 688,368,653 688,491,533 688,491,546
($\pm 0.0$) ($\pm 2.75\times 10^{-1}$) ($\pm 6.05\times 10^{-2}$) ($\pm 4.04\times 10^{-2}$) ($\pm 0.0$) ($\pm 1.11\times 10^{-1}$) ($\pm 3.20\times 10^{-2}$) ($\pm 3.49\times 10^{-2}$)
15 $1.14\times 10^{1}$ $8.76\times 10^{-1}$ $4.65\times 10^{-1}$ $6.35\times 10^{-1}$ $1.23\times 10^{1}$ $5.47\times 10^{0}$ $5.08\times 10^{0}$ $5.20\times 10^{0}$ 184,320 688,389,138 688,573,458 688,573,476
($\pm 0.0$) ($\pm 4.36\times 10^{-2}$) ($\pm 2.43\times 10^{-2}$) ($\pm 3.55\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.02\times 10^{-1}$) ($\pm 3.22\times 10^{-2}$) ($\pm 3.59\times 10^{-2}$)
20 $9.71\times 10^{0}$ $8.96\times 10^{-1}$ $3.07\times 10^{-1}$ $5.06\times 10^{-1}$ $1.09\times 10^{1}$ $5.49\times 10^{0}$ $4.85\times 10^{0}$ $5.21\times 10^{0}$ 245,760 688,409,623 688,655,383 688,655,406
($\pm 0.00$) ($\pm 5.07\times 10^{-2}$) ($\pm 1.90\times 10^{-2}$) ($\pm 2.83\times 10^{-2}$) ($\pm 0.00$) ($\pm 6.92\times 10^{-2}$) ($\pm 3.01\times 10^{-2}$) ($\pm 3.14\times 10^{-2}$)
$32^{3}$ 5 $2.02\times 10^{1}$ $9.15\times 10^{0}$ $1.84\times 10^{1}$ $5.56\times 10^{0}$ $2.04\times 10^{1}$ $1.05\times 10^{1}$ $1.87\times 10^{1}$ $7.16\times 10^{0}$ 491,520 688,505,864 688,997,384 688,997,392
($\pm 0.00$) ($\pm 1.10\times 10^{1}$) ($\pm 2.61\times 10^{-1}$) ($\pm 1.03\times 10^{-1}$) ($\pm 0.00$) ($\pm 1.05\times 10^{1}$) ($\pm 2.34\times 10^{-1}$) ($\pm 7.28\times 10^{-2}$)
10 $1.42\times 10^{1}$ $1.52\times 10^{1}$ $1.23\times 10^{1}$ $3.98\times 10^{0}$ $1.47\times 10^{1}$ $1.63\times 10^{1}$ $1.32\times 10^{1}$ $6.50\times 10^{0}$ 983,040 688,669,709 689,652,749 689,652,762
($\pm 0.00$) ($\pm 1.74\times 10^{1}$) ($\pm 1.45\times 10^{-1}$) ($\pm 6.67\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.66\times 10^{1}$) ($\pm 1.27\times 10^{-1}$) ($\pm 5.38\times 10^{-2}$)
15 $1.15\times 10^{1}$ $1.46\times 10^{1}$ $9.18\times 10^{0}$ $2.90\times 10^{0}$ $1.22\times 10^{1}$ $1.59\times 10^{1}$ $1.06\times 10^{1}$ $6.18\times 10^{0}$ 1,474,560 688,833,554 690,308,114 690,308,132
($\pm 0.00$) ($\pm 1.76\times 10^{1}$) ($\pm 1.81\times 10^{-1}$) ($\pm 5.54\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.68\times 10^{1}$) ($\pm 1.55\times 10^{-1}$) ($\pm 4.65\times 10^{-2}$)
20 $9.78\times 10^{0}$ $1.10\times 10^{1}$ $6.91\times 10^{0}$ $2.21\times 10^{0}$ $1.08\times 10^{1}$ $1.26\times 10^{1}$ $8.86\times 10^{0}$ $5.97\times 10^{0}$ 1,966,080 688,997,399 690,963,479 690,963,502
($\pm 0.00$) ($\pm 1.52\times 10^{1}$) ($\pm 2.84\times 10^{-1}$) ($\pm 4.70\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.44\times 10^{1}$) ($\pm 1.55\times 10^{-1}$) ($\pm 3.96\times 10^{-2}$)
$64^{3}$ 5 $2.02\times 10^{1}$ $2.89\times 10^{1}$ $1.95\times 10^{1}$ $9.00\times 10^{0}$ $2.04\times 10^{1}$ $2.89\times 10^{1}$ $1.96\times 10^{1}$ $9.53\times 10^{0}$ 3,932,160 689,767,432 693,699,592 693,699,600
($\pm 0.00$) ($\pm 1.94\times 10^{1}$) ($\pm 1.53\times 10^{-1}$) ($\pm 9.62\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.91\times 10^{1}$) ($\pm 1.48\times 10^{-1}$) ($\pm 9.27\times 10^{-2}$)
10 $1.42\times 10^{1}$ $2.50\times 10^{1}$ $1.35\times 10^{1}$ $7.01\times 10^{0}$ $1.46\times 10^{1}$ $2.52\times 10^{1}$ $1.40\times 10^{1}$ $7.95\times 10^{0}$ 7,864,320 691,078,157 698,942,477 698,942,490
($\pm 0.00$) ($\pm 2.00\times 10^{1}$) ($\pm 1.22\times 10^{-1}$) ($\pm 7.07\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.96\times 10^{1}$) ($\pm 1.18\times 10^{-1}$) ($\pm 6.18\times 10^{-2}$)
15 $1.15\times 10^{1}$ $1.65\times 10^{1}$ $1.08\times 10^{1}$ $5.59\times 10^{0}$ $1.22\times 10^{1}$ $69\times 10^{1} 1.1$$16\times 10^{1}$ $7.01\times 10^{0}$ 11,796,480 692,388,882 704,185,362 704,185,380
($\pm 0.00$) ($\pm 1.65\times 10^{1}$) ($\pm 1.79\times 10^{-1}$) ($\pm 3.82\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.62\times 10^{1}$) ($\pm 1.69\times 10^{-1}$) ($\pm 3.63\times 10^{-2}$)
20 $9.78\times 10^{0}$ $1.94\times 10^{1}$ $9.13\times 10^{0}$ $4.54\times 10^{0}$ $1.08\times 10^{1}$ $1.97\times 10^{1}$ $1.02\times 10^{1}$ $6.45\times 10^{0}$ 15,728,640 693,699,607 709,428,247 709,428,270
($\pm 0.00$) ($\pm 1.82\times 10^{1}$) ($\pm 9.70\times 10^{-2}$) ($\pm 3.73\times 10^{-2}$) ($\pm 0.00$) ($\pm 1.78\times 10^{1}$) ($\pm 9.29\times 10^{-2}$) ($\pm 3.39\times 10^{-2}$)